### v1

# Quantum Emission from Coupled Spin Pairs in Hexagonal Boron Nitride

Abstract
Optically addressable defect qubits in wide band gap materials are favorable candidates for room temperature quantum information processing. The twodimensional (2D) hexagonal boron nitride (hBN) is an attractive solid state platform with a great potential for hosting bright quantum emitters with quantum memories with leveraging the potential of 2D materials for realizing scalable preparation of defect qubits. Although, room temperature bright defect qubits have been recently reported in hBN but their microscopic origin, the nature of the optical transition as well as the optically detected magnetic resonance (ODMR) have been remained elusive. Here we connect the photostability and spectral diffusion of quantum emitters to donor-acceptor pairs (DAP) in hBN by means of ab initio calculations. We find that DAPs can exhibit ODMR signal for the acceptor counterpart of the defect pair with $S=1/2$ ground state at non-zero magnetic fields depending on the donor partner. The donor-acceptor pair model and its transition mechanisms provide a recipe towards defect qubit identification and performance optimization in hBN for quantum applications.
I Introduction
Optically active atomic defects in wide band gap materials serve as promising single photon emitters (SPE) which are key requirements for quantum information technologies Maze et al. (2008); Gruber et al. (1997); Balasubramanian et al. (2009); Togan et al. (2010); Weber et al. (2010); Wolfowicz et al. (2021); Chatterjee et al. (2021); Bassett et al. (2019). Hexagonal boron nitride (hBN) is a layered van der Waals (vdW) material and a favorable SPE host due to its versatility of fabrication and compatibility with lithographic processing Cassabois et al. (2016); Caldwell et al. (2019); Sajid et al. (2020); Liu et al. (2022a); Montblanch et al. (2023). The measured single photon emission in hBN features high brightness, room temperature stability, sharp zero-phonon-lines (ZPL) and short excited state lifetime in visible wavelength region Tran et al. (2016a, b); Stern et al. (2022); Mendelson et al. (2021); Guo et al. (2023); Noh et al. (2018); Xu et al. (2021); Kianinia et al. (2018); Li et al. (2019). These exceptional properties motivate researchers to utilize the spin defect in hBN for quantum sensing and imaging Gao et al. (2023a); Kumar et al. (2022); Lyu et al. (2022); Gao et al. (2023b); Yang et al. (2022); Gottscholl et al. (2021a); Xiong et al. (2024). The emissions observed in experiment are classified into two groups based on their ZPL (Class-I and II) and phonon sideband (PSB) spectrum Tran et al. (2016b). The difference indicates that the emissions might come from various defects, and the local environment also has significant influence Li et al. (2024); Plo et al. (2024); Iwański et al. (2024). However, one major challenge is the identification of exact defect structures, which is the prerequisite for realizing deterministic formation and control. Although numerous defects are proposed theoretically with parameters that can match experimental result, direct one-to-one identification through electron microscopy is imperative.
Boron vacancy (VB), which has been unambiguously identified, is a native defect in hBN and received intense research effort Jin et al. (2009); Abdi et al. (2018); Gottscholl et al. (2020); Ivády et al. (2020); Gottscholl et al. (2021b); Liu et al. (2022b); Murzakhanov et al. (2022); Haykal et al. (2022); Mu et al. (2022); Durand et al. (2023). The negatively charged VB has a triplet ground state manifesting its potential in quantum sensing operating at room temperature, although, it is an intrinsically dim emitter Libbi et al. (2022) which makes it difficult to use it at single spin level. Beside VB, the coherent control of single electron spin has also been realized on other color centers Chejanovsky et al. (2021), some of them operating at room temperature Horder et al. (2022); Stern et al. (2024). However, the origin of these color centers has remained unidentified. Carbon and oxygen are common impurities in hBN Maity et al. (2021); Watanabe et al. (2004) that may constitute the color centers. Several candidates with vacancies are proposed such as (CNVB and VNCBTran et al. (2016b); Tawfik et al. (2017); Sajid et al. (2018) and (ONVB and VNOBNoh et al. (2018); Li and Gali (2022). In addition, substitutional carbon defects are believed to highly involved during the formation of the color centers and carbon clusters with various size could have emission not only at visible range Chejanovsky et al. (2021); Mendelson et al. (2021); Li et al. (2022a) but also ultraviolet region as observed in hBN sample Bourrellier et al. (2016); Vuong et al. (2016); Mackoit-Sinkevičienė et al. (2019); Li et al. (2022b).
Electron paramagnetic resonance (EPR) is a powerful tool to identify paramagnetic defects in semiconductors. Earlier work reported three-boron center (TBC) and the one boron center (OBC) with resolvable hyperfine splitting from boron nuclei Katzir et al. (1975); Andrei et al. (1976). Recent experiment observed $S=1/2$ paramagnetic defect with strong hyperfine interaction of two equivalent nitrogen nuclei Toledo and Krambrock (2020) and we assigned this signal to the negatively charged ONVB defect based on the excellent agreement between the experimental and simulated EPR spectra Li and Gali (2022). Noteworthy, the existence of ONVB defect has been confirmed by a following annular dark-field scanning transmission electron microscopy (ADF-STEM) experiment Li et al. (2023). In addition, carbon and oxygen substitutions are observed simultaneously nearby with the same technique. This is a strong evidence that an extra charge on ONVB defect giving rise to the EPR signal could come from the carbon and oxygen substitution of boron (CB) and nitrogen (ON), respectively, with donor character Weston et al. (2018). In other words, CB or ON forms a donor-acceptor pair (DAP) with ONVB that may be described as CB+ – ONVB- or ON+ – ONVB- in the ground state which has $S=1/2$ ground state coming from the spin density around the ONVB- part of the DAP.
Here we perform comprehensive theoretical calculations for the optical properties of DAP with different separation distance. We find with our model that the donor (CB and ON) indeed provides an electron towards ONVB defect which becomes negatively charged. The distance between donor and acceptor can significantly influence the electronic structure. This could be a possible explanation for the ZPL variation observed in experiments. We find that ON – ONVB DAP is photostable with quantum yields similar to that of the isolated negatively charged ONVB. On the contrary, the recombination energy of CB – ONVB develops metastable dim states which act as non-radiative decay pathways. The spin flipping in these metastable dim states could mix the doublet and quartet multiplets which causes spin polarization at the $S=1/2$ ground state of the negatively charged ONVB when external magnetic field splits these Kramers-doublets. As a consequence, ODMR spectrum Li and Gali (2022) may be observed for the pair of CB – ONVB in the $S=1/2$ ground state in the presence of a constant magnetic field.
II Results
Figure 1: (a) Schematics view of CB – ONVB DAP with various distances. The numbers labeled here indicate the location of CB. (b) The distance-dependent ZPL values. (c) The electronic structure of defects in the ground state. Number 10 is the isolated ONVB(--) defect for reference.
ADF-STEM technique is capable of imaging carbon and oxygen atoms near the VB monovacancy, particularly focusing on the innermost nitrogen atoms Li et al. (2023). However, these monovacancy structures with first-neighbor atom substitutions do not yield the bright emission at 2 eV that is frequently observed in experiments, as discussed in Supplementary Note 1. Instead, another interesting phenomenon observed experimentally involves the carbon and oxygen substitutional defects in vacancy-free regions. The presence of CB, CN and ON can be readily confirmed, while OB has a high formation energy and therefore is rarely observed Li et al. (2023). A previous theoretical study has indicated the donor characteristics of CB and ON Weston et al. (2018). We emphasize that the oxygen related defect ONVB may exhibit a 2 eV emission in its negative charge state Li and Gali (2022). Hence, we propose that the observed SPE is ONVB-, with the additional charge supplied by either CB or ON. In Supplementary Figure 1, we illustrate that the single electron levels of CB or ON are situated at higher energies than the empty state $a_{1}$ of ONVB  promoting electron transfer from the donors to ONVB. Furthermore, the total energy of CB+ – ONVB- configuration is lower than that of CB0 – ONVB0, confirming the stability of DAP. Noteworthy, the charge transfer within DAP is usually distance-dependent; in wide band gap materials the probability of charge transfer rapidly decays with increasing the distance between the donor and acceptor. Therefore, we focus on several DAP structures characterised by relatively short distances.
We begin by examining the properties of the CB – ONVB DAP, as illustrated in Fig. 1. The first-neighbor site, which is capable of forming a direct bond to oxygen, was not considered due to a potential repulsion interaction between ON and CB Guo et al. (2023). CB1 and CB2 sites are close to the oxygen atom, leading to the remarkable influence of the donor on the properties of the DAPs. Following the previously defined notations, as shown in Supplementary Figure 1, the optical transition within ONVB- occurs between the $a_{1}$ and $b_{1}$ levels. However, the strong interaction within CB1 and CB2 can change the order of occupied defect levels of ONVB- and the $a_{1}$ level shifts down while $a_{2}$ or $b_{2}$ level shifts up (the presence of CB makes the symmetry $C_{1h}$ so the original $a_{2}$ and $b_{2}$ labels are unified). Since the $a_{2}$ ($b_{2}$) orbital has out-of-plane wavefunctions while $b_{1}$ lies in-plane, the transition dipole moment is along out-of-plane direction with very small magnitude at $\sim$0.12 Debye so the optical transition between them is dim. The bright emission still comes from the $a_{1}$ to $b_{1}$. The large energy separation between $a_{1}$ and $b_{1}$ leads large ZPL, it is 3.7 and 3.9 eV for CB1 and CB2, respectively, and locates at ultraviolet region. With increased distance, the energy levels’ order restores to the ONVB- character and the wavefunction of CB retrieves $D_{3h}$ symmetry. Another trend is the unoccupied levels of CB shift downward while the $b_{1}$ level shifts upward. In CB5 case, the empty level from CB is below the $b_{1}$ orbital from ONVB. Nevertheless, the transition dipole moment from $a_{1}$ to CB is only 0.16 Debye so we focus on the bright emission localized on ONVB where the ZPL gradually converges to 1.97 eV.
The above discussed bright emission is a local excitation within ONVB-. Beside this, the charge transfer process from ONVB- to CB+ may also luminesce, although it is dim. The dim metastable state (MS) has two sub-states $S=1/2$ doublet or $S=3/2$ quartet depending on relative spin orientations within DAP. With Kohn-Sham density functional theory, we could reliably calculate the total energy and geometry of the $S=3/2$ state. Supplementary Figure 3 shows the Kohn-Sham energy difference of the $S=1/2$ dim and bright transitions. Generally the bright local excitation is independent on the distance between DAP while the dim one shows strong relation. The bright transition is the first optical transition in CB3 and CB4 and when CB resides at larger distances from ONVB then the dim transition has lower energy. This means there could be a observable non-radiative decay path for the local excited state to decay to the ground state. A simple model to capture the dim transition mechanism can be expressed as
$E(R_{i})=E_{\text{gap}}-(E_{\text{D}}+E_{\text{A}})+\frac{e^{2}}{\epsilon R_{i}}\text{,}$ (1)
where $E_{\text{D}}$ and $E_{\text{A}}$ are the energy levels of the donor and the acceptor in the band gap, respectively. Here, we use the charge transition level (CTL) of CB($0|+1$) and ONVB($-1|0$) from previous study. $\epsilon$ is the dielectric constant of hBN and $R_{i}$ is separated distance between donor and acceptor. The last term is the Coulombic interaction and the reciprocal function leads fast convergence of $E(R_{i})$ to the CTL difference (1.27 eV) of DAP with large $R_{i}$, as shown in Supplementary Figure 4. The $S=3/2$ state can converge properly and it might be a reliable reference to predict the energy of $S=1/2$ state because the energy difference here is from the spin-spin exchange interaction which is relatively small. The total energy of $S=3/2$ state lies with 1.68 eV to 2.36 eV higher than that of the ground state depending on the DAP distance and this means that the bright excited state ($\sim$2.0 eV) and the metastable states can switch their energy position.
ON – ONVB is different from CB – ONVB case. Even with short distance between them, the ONVB- character is still well kept, and this might be due to shallow donor property of ON. The empty levels from ON are very close to CBM so they have negligible influence on the optical transition within ONVB. Except the ON1 site, the ZPL energies are all below 2.23 eV. Fig. 2a shows the transition dipole moment evolution upon the configurations considered. It converges fast to that of the isolated ONVB- with slightly increased distance between DAP. The large energy difference between ON($0|+1$) and ONVB($-1|0$) results in $E(R_{i})$ to be higher than 2.89 eV, e.g., the $S=3/2$ level of ON4 lies at 3.59 eV above the ground state. Correspondingly, no metastable states arise for ON+ – ONVB-, which makes the optically and spin active ONVB- counterpart of DAP a photostable emitter.
Figure 2: (a) Schematics view of ON – ONVB DAP with various distances. The numbers labeled here indicate the location of ON. (b) The distance-dependent ZPL values. (c) The electronic structure of defects in the ground state. Number 10 is the isolated ONVB- defect for reference.
Figure 3: (a) The transition dipole moment evolution upon defect configurations we considered. The dash line indicates the transition dipole moment of ONVB-. (b)(c) Mechanism of DAP transition in semiconductors.
III Discussion
The fluorescence mechanisms within DAP as shown in Fig. 3b,c can be classified in two types according to the relative energy levels of the donor and the acceptor. If the donor is lower than the acceptor level then the charge transfer cannot occur spontaneously. Charge transfer can be induced by illumination from $D^{0}$ to $A^{0}$ (neutral charge state) with photon energy $h\nu_{a}$, and this creates ionized donor $D^{+}$ and acceptor $A^{-}$. In other words, the electron from occupied state of $D^{0}$ is excited to the empty state of $A^{0}$. The electron goes back through ensuing radiative decay (charge recombination) and one photon is emitted with energy $h\nu_{e}$. If the donor level is higher than the acceptor level then the electron from donor may spontaneously transfer to the acceptor when they are located in short distance. Then the photoexcitation converts the charged DAP to neutral. The charge transfer from donor to acceptor is distance-dependent and a recent study investigated the charge transfer dynamics based on Marcus theory framework Lee et al. (2019). Unlike the above fluorescence mechanisms within DAP, here the photoexcitation is generally active on the orbitals localized on the acceptor, which has a bright optical transition, especially for ON+ – ONVB-. The shallow donor makes the levels of the dim ON0 – ONVB0 states to lie at much higher energies than the bright excited state’s level. In CB+ – ONVB- DAP the situation is different, as its optical loop depicted in Fig. 4. The optical excitation first drives the doublet ground state to the bright excited state. The small energy difference between bright excited state and the dim metastable states enables the charge transfer from ONVB to CB through non-radiative decay. The doublet ($D_{s}$) metastable level and the quartet ($Q_{s}$) metastable level are separated by the spin-spin exchange interaction $J$. The ground state and the bright doublet excited states are linked to the $D_{s}$ state by either dim optical transition or internal conversion (IC) while they are linked to the $Q_{s}$ by intersystem crossing (ISC) where spin-flip mediates the latter. By taking the same orbitals in the $D_{s}$ and $Q_{s}$ it is a common wisdom that IC is a much faster process than ISC. Therefore, the $D_{s}$ state will be populated in the metastable manifolds via IC. At zero magnetic field the $D_{s}$ state provides radiative and non-radiative pathways which shorten the optical lifetime of the ONVB- in this DAP when compared to the optical lifetime of ON+ – ONVB- DAP.
At non-zero magnetic fields the Kramers-doublets split in CB+ – ONVB- DAP where the presence of $Q_{s}$ plays an important role in the spin-dependent non-radiative decay. For instance, the higher branch $m_{S}=+1/2$ in $D_{s}$ can mix with lower branch of $m_{S}=-3/2$ in $Q_{s}$. In other word, the $D_{s}$ acquires some $Q_{s}$ character and vice versa. Therefore, the $m_{S}=-1/2$ spin sublevel localized on ONVB- in $D_{s}$ can have larger population than $m_{S}=+1/2$ through the non-radiative decay $r3$ (see Fig. 4). This will populate the $m_{S}=-1/2$ spin state over the $m_{S}=+1/2$ spin state in the electronic ground state which makes the $m_{S}=-1/2$ spin level brighter than that of $m_{S}=+1/2$. As a consequence, ODMR signal can be observed for an $S=1/2$ ground state. The key characteristic here is the quasi-degenerate $D_{s}$ and $Q_{s}$ states in the metastable CB0 – ONVB0 configuration which is a spin pair of $S=1/2$ (see Ref. Auburger and Gali (2021)) and $S=1$ defects (see Ref. Li and Gali (2022)), respectively. This can lead to spin-polarization of a split $S=1/2$ Kramers ground state doublet as observed in ODMR.
Actually, such effects for DAP systems have been already reported in wide band gap materials such as zinc-sulfide Nicholls et al. (1979) and diamond Nazare et al. (1995) but only phenomenological models have been provided to understand this effect. The intersystem crossing processes for spin pairs have been mostly discussed for organic chromophores Ishii et al. (1998); Kandrashkin and van der Est (2003); Kandrashkin and Van Der Est (2004); Teki (2020); Qiu et al. (2023) which includes dipolar mixing ($\hat{H}_{\text{di}}$), spin-orbit-coupling ($\hat{H}_{\text{soc}}$) and hyperfine couplings ($\hat{H}_{\text{hf}}$). However, it should be noted that none of these can be easily determined due to the distance-dependence and $D_{s}$-$Q_{s}$ splitting-dependence. The $D_{s}$-$Q_{s}$ splitting scales with distance as $R^{-6}$. Relatively, the intersystem crossing caused by hyperfine interaction decays slower than the dipolar mixing and spin-orbit-coupling. However, this theory might fail with large separation distance between the two defects, e.g., $\sim 5$ nm, when the wavefunction overlap is negligible. For specified distance $R_{i}$, the rate to flip can be evaluated by Fermi Golden rule as
$\Gamma=\frac{2\pi}{\hbar}|\langle\psi(D_{s})|\hat{H}_{\text{di}}+\hat{H}_{\text{soc}}+\hat{H}_{\text{hf}}|\psi(Q_{s})\rangle|^{2}L(J)\text{,}$ (2)
where $\psi(D_{s})$ and $\psi(Q_{s})$ are the doublet and quartet wavefunction, and $L(J)$ is the lineshape function associated with the phonon overlaps between the two states. With the strongest hyperfine coupling of 55 MHz from the two 14N nuclear spins of ONVB in CB5, the estimated rate has upper bound at $10^{8}$ Hz when we assume the doublet and quartet have nearly identical energy and geometry. This upper bound might be suppressed, e.g., via $L(J)$. Nevertheless, it still represents a sufficiently fast process for an efficient spin mixing.
The distance between the DAP constituents should be long enough to reduce the energy gap $J$ between $D_{s}$ and $Q_{s}$ but not too long to minimize the spin-mixing rate. Unfortunately, the direct calculation of $D_{s}$ and $J$ is beyond the scope of the present study because it requires multireference methods that cannot be applied to the supercell size required for the DAP model. Nevertheless, an effective Hamiltonian for ONVB0 – CB0 with $S_{1}=1$ and $S_{2}=1/2$ coupled spin pairs at a given magnetic field $B$ may be written as
$H_{\text{spin}}=S_{1}\mathbf{D}S_{1}+\sum_{i}H^{(i)}_{\text{hf}}+gBS_{1}+gBS_{2}+JS_{1}S_{2}\text{,}$ (3)
where the gyromagnetic factor of the electrons is $g\approx 2.0023$ (very low spin-orbit interaction) for each constituent, $J$ is the isotropic exchange interaction between the spin pairs, $\mathbf{D}$ is the $D$-tensor of ONVB0 [$D=3.8$ GHz and $E=0.091$ GHz], and we only use the largest hyperfine couplings of the two 14N $I=1$ spins. We numerically solve Eq. (3) as a function of $B$-field (set perpendicular to the hBN sheet) and the $J$ spin exchange interaction as implemented in Easyspin Stoll and Schweiger (2006).
Figure 4: The optical pumping loop of CB+ ONVB. (a) Bright excitation (red) from ground state (GS) to excited state (ES) is localized on ONVB(--). Through internal conversion, the ES state can go to doublet IS CB(0) + ONVB(0). The doublet IS can transfer to quartet IS through intersystem crossing with energy separation $J>0$. The $\pm$ 1/2 and 3/2 spin sublevels are separated by zero field splitting in quartet. The black dot indicates spin mixing between doublet and quartet. The black circle indicate the population magnitude at ground state. (b) The spin sublevel splitting under magnetic field in doublet and quartet. The $J$ is --3000, 0 and 3000 MHz, accordingly. (c) The population ratio between $|D_{s},+1/2\rangle$ and $|D_{s},-1/2\rangle$ at dim state upon magnetic field.
Fig. 4b shows the splitting of spin sublevels under magnetic field for various $J$, which can exhibit different magnitude and sign depending on the relative orientation of coupled spins as appears in our DFT calculation. Due to the zero-field splitting of triplet sublevels of ONVB  the $|Q_{s},\pm 1/2\rangle$ always lie low and the main spin mixing is from $|Q_{s},\pm 3/2\rangle$ and $|D_{s},\pm 1/2\rangle$. The mixing is largely due to $|D_{s},-1/2\rangle$ and $|Q_{s},+3/2\rangle$ when $J<0$, while the mixing is between $|D_{s},+1/2\rangle$ and $|Q_{s},-3/2\rangle$ when $J>0$. The different coefficient of spin sublevels in eigenstates indicates the distinguishable population and this population difference, or ODMR contrast, could be tuned through magnetic field. More specifically, the variation in mixing with the quartet states can result in different non-radiative relaxation rates, leading to hyperpolarization and ODMR contrast. For each finite magnitude of J-coupling, we observe a strong dependence of the ratio of mixing coefficients on the magnetic field, as illustrated in Fig. 4b. This difference is particularly pronounced when the applied field is set close to the avoided crossing, reaching a maximum value of approximately 1.4. Ultimately, this should translate into differences in the internal conversion rates to the ground state. It is important to note that for $J$=0, i.e., when two non-interacting defects are present, both doublet states mix equally with the quartet manifold at each magnetic field, resulting in no discernible difference in non-radiative decay.
To further substantiate the role of spin mixing in mediating the ODMR signal, we calculated the internal conversion rates from the metastable state. We approach this process by using the excited state and MS quartet for $r1$, MS quartet and ground state for $r2$, separately. With short distance below 1 nm, the small energy difference between bright and dim state and strong electron-phonon coupling lead to fast internal conversion rate around $10^{10}$ MHz. Therefore the system relax back to ground state rapidly without detectable optical signal. The orbital overlap integral and electron-phonon coupling decay with increasing distance where the $R$ distant dependent value of the integrals may be approximated by the overlap of two Slater orbitals which goes with $e^{-R}$ Silver and Ruedenberg (1968). By the two configurations CB5 and CB7, we can extrapolate the electron-phonon coupling and conversion rate to longer distance that cannot be simulated with DFT, as discussed in Supplementary Note 2. The estimated conversion rate $r1$ can be 6-8 MHz with 18 Å separation.
With around 2 nm separation, the present coupled spin pair model could explain the previously reported signature of the ODMR signal from ground state $S=1/2$ defects reported in hBN sheets and nanotubes Scholten et al. (2023); Gao et al. (2023c); Guo et al. (2023). We note that if the metastable state is sufficiently long-living to carry out spin operations (e.g., ST1 defect in diamond Lee et al. (2013)) then this spin pair model can be extended to the ODMR centers where the spin-polarization is observed in the metastable state. In this case, the ground state could be singlet from an $S=1/2$ acceptor and $S=1/2$ donor whereas the metastable states comprise single and triplet spin pairs. We note a recent work (Ref. Scholten et al. (2024)) during the preparation of our manuscript which observes this phenomenon without adding microscopic models for the underlying processes.
In summary, we propose the DAP emission is responsible for most of the visible ZPL emission at around 2.0 eV. The variance in the ZPL wavelengths and optical lifetimes reported in the experiments can be explained by the type and the distance of the donor defect with respect to a key acceptor defect, ONVB. Our simulations reveal the ODMR mechanism associated with the spin pairs in certain DAP structures activated in the dim metastable states which is manifested as ODMR detection of the $S=1/2$ ground state at non-zero constant magnetic fields. DAPs in wide band gap semiconductors and materials are quite common structures and are responsible for optical emissions and ODMR signal for Kramers doublet spin systems. The general mechanisms proposed here is insightful to not only the defects engineering in hBN for quantum information processing but also the optoelectronic applications of defects in other semiconductors.
IV Methods
In this paper, we carry out the density functional theory (DFT) calculations by the Vienna ab initio simulation package (VASP) code Kresse and Furthmüller (1996a, b) with plane wave basis set. The valence electrons and core part are described with projector augmented wave (PAW) potentials Blöchl (1994); Kresse and Joubert (1999). A $8\times 8$ two-layered supercell model was used to avoid the interactions between periodic images and this is sufficient to use the $\Gamma$-point sampling scheme. The interlayer vdW interaction was described with DFT-D3 method of Grimme  Grimme et al. (2010). With mixing parameter $\alpha=0.32$, the hybrid density functional of Heyd, Scuseria, and Ernzerhof (HSE)  Heyd et al. (2003) could yield experimental optical gap around 6 eV without renormalization energy included and this functional is used to optimize the geometry and calculate electronic structures. The convergence threshold for the forces was set to 0.01 eV/Å with cutoff energy at 450 eV. $\Delta$SCF method Gali et al. (2009) was used to calculate excited states. The band alignment is based on the core level energy of one atom far away from the defect center in defective and it in perfect supercell.
Author contribution
S. L. and A. G. designed the project. All authors discussed the results and contributed to the manuscript writing.
Competing interests
The authors declare that there are no competing interests.
Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Acknowledgements.
Support by the Ministry of Culture and Innovation and the National Research, Development and Innovation Office within the Quantum Information National Laboratory of Hungary (Grant No. 2022-2.1.1-NL-2022-00004) as well as the European Commission for the projects QuMicro (Grant No. 101046911) and SPINUS (Grant No. 101135699) are much appreciated. AG acknowledges the high-performance computational resources provided by KIFÜ (Governmental Agency for IT Development of Hungary).

### v2

@@
 Optically addressable defect qubits in wide band gap materials are favorable candidates for room temperature quantum information processing.
-The twodimensional (2D) hexagonal boron nitride (hBN) is an attractive solid state platform with a great potential for hosting bright quantum emitters with quantum memories with leveraging the potential of 2D materials for realizing scalable preparation of defect qubits.
+The two-dimensional (2D) hexagonal boron nitride (hBN) is an attractive solid state platform with a great potential for hosting bright quantum emitters with quantum memories with leveraging the potential of 2D materials for realizing scalable preparation of defect qubits.
 Although, room temperature bright defect qubits have been recently reported in hBN but their microscopic origin, the nature of the optical transition as well as the optically detected magnetic resonance (ODMR) have been remained elusive.
-Here we connect the photostability and spectral diffusion of quantum emitters to donor-acceptor pairs (DAP) in hBN by means of ab initio calculations.
+Here we connect the variance in the optical spectra, optical lifetimes and spectral stability of quantum emitters to donor-acceptor pairs (DAP) in hBN by means of ab initio calculations.
 We find that DAPs can exhibit ODMR signal for the acceptor counterpart of the defect pair with $S=1/2$ ground state at non-zero magnetic fields depending on the donor partner.
@@
 I Introduction
-Optically active atomic defects in wide band gap materials serve as promising single photon emitters (SPE) which are key requirements for quantum information technologies Maze et al.
-(2008); Gruber et al.
-(1997); Balasubramanian et al.
-(2009); Togan et al.
-(2010); Weber et al.
+Isolated optically active atomic defects in wide band gap materials serve as single photon emitters (SPE) which are key requirements for quantum information technologies Zhang et al.
 (2020); Wolfowicz et al.
-(2021); Chatterjee et al.
-(2021); Bassett et al.
 (2021).
@@
 (2020); Liu et al.
-(2022a); Montblanch et al.
+(2022); Montblanch et al.
 (2023).
-The measured single photon emission in hBN features high brightness, room temperature stability, sharp zero-phonon-lines (ZPL) and short excited state lifetime in visible wavelength region Tran et al.
+The observed SPEs in hBN feature high brightness, room temperature stability, sharp zero-phonon-lines (ZPL) peaks at around 2 eV and short excited state lifetime Tran et al.
 (2016a, b); Stern et al.
@@
 (2019).
-These exceptional properties motivate researchers to utilize the spin defect in hBN for quantum sensing and imaging Gao et al.
-(2023a); Kumar et al.
-(2022); Lyu et al.
-(2022); Gao et al.
-(2023b); Yang et al.
-(2022); Gottscholl et al.
-(2021a); Xiong et al.
-(2024).
-The emissions observed in experiment are classified into two groups based on their ZPL (Class-I and II) and phonon sideband (PSB) spectrum Tran et al.
+Furthermore, the coherent control of single electron spins has been recently realized in hBN Chejanovsky et al.
+(2021), some of them operating at room temperature Horder et al.
+(2022); Stern et al.
+(2024) where the electron spin initialization and readout is based on optical excitation and emission of the defect spins, called optically detected magnetic resonance (ODMR).
+One major challenge is the identification of exact defect structures of the SPEs and single-spin ODMR centers which is a prerequisite for realizing deterministic formation and control.
+The observed photoluminescence (PL) spectra exhibit various ZPL energies and phonon sideband (PSB) in the spectrum and many of them show similar optical lineshapes Tran et al.
 (2016b).
-The difference indicates that the emissions might come from various defects, and the local environment also has significant influence Li et al.
+The observed emissions might come from various defects but the similarities in their optical lineshape also imply common defect types in diverse crystalline environment Li et al.
 (2024); Plo et al.
@@
 (2024).
-However, one major challenge is the identification of exact defect structures, which is the prerequisite for realizing deterministic formation and control.
-Although numerous defects are proposed theoretically with parameters that can match experimental result, direct one-to-one identification through electron microscopy is imperative.
-Boron vacancy (VB), which has been unambiguously identified, is a native defect in hBN and received intense research effort Jin et al.
-(2009); Abdi et al.
-(2018); Gottscholl et al.
-(2020); Ivády et al.
-(2020); Gottscholl et al.
-(2021b); Liu et al.
-(2022b); Murzakhanov et al.
-(2022); Haykal et al.
-(2022); Mu et al.
-(2022); Durand et al.
-(2023).
-The negatively charged VB has a triplet ground state manifesting its potential in quantum sensing operating at room temperature, although, it is an intrinsically dim emitter Libbi et al.
-(2022) which makes it difficult to use it at single spin level.
-Beside VB, the coherent control of single electron spin has also been realized on other color centers Chejanovsky et al.
-(2021), some of them operating at room temperature Horder et al.
-(2022); Stern et al.
-(2024).
-However, the origin of these color centers has remained unidentified.
-Carbon and oxygen are common impurities in hBN Maity et al.
-(2021); Watanabe et al.
-(2004) that may constitute the color centers.
-Several candidates with vacancies are proposed such as (CNVB and VNCBTran et al.
-(2016b); Tawfik et al.
-(2017); Sajid et al.
-(2018) and (ONVB and VNOBNoh et al.
-(2018); Li and Gali (2022).
-In addition, substitutional carbon defects are believed to highly involved during the formation of the color centers and carbon clusters with various size could have emission not only at visible range Chejanovsky et al.
-(2021); Mendelson et al.
-(2021); Li et al.
-(2022a) but also ultraviolet region as observed in hBN sample Bourrellier et al.
-(2016); Vuong et al.
-(2016); Mackoit-Sinkevičienė et al.
-(2019); Li et al.
-(2022b).
-Electron paramagnetic resonance (EPR) is a powerful tool to identify paramagnetic defects in semiconductors.
-Earlier work reported three-boron center (TBC) and the one boron center (OBC) with resolvable hyperfine splitting from boron nuclei Katzir et al.
-(1975); Andrei et al.
-(1976).
-Recent experiment observed $S=1/2$ paramagnetic defect with strong hyperfine interaction of two equivalent nitrogen nuclei Toledo and Krambrock (2020) and we assigned this signal to the negatively charged ONVB defect based on the excellent agreement between the experimental and simulated EPR spectra Li and Gali (2022).
+An $S=1/2$ paramagnetic defect with strong hyperfine interaction of two equivalent nitrogen nuclei has been observed by electron paramagnetic resonance (EPR) Toledo and Krambrock (2020) and we assigned this signal to the negatively charged ONVB defect, i.e., oxygen substituting nitrogen adjacent to boron-vacancy, based on the excellent agreement between the experimental and simulated EPR spectra Li and Gali (2022).
 Noteworthy, the existence of ONVB defect has been confirmed by a following annular dark-field scanning transmission electron microscopy (ADF-STEM) experiment Li et al.
@@
 In other words, CB or ON forms a donor-acceptor pair (DAP) with ONVB that may be described as CB+ – ONVB- or ON+ – ONVB- in the ground state which has $S=1/2$ ground state coming from the spin density around the ONVB- part of the DAP.
+In this sense, the common defect type is ONVB acceptor defect and the variance in the optical properties comes from the type and location of the donor partner.
 Here we perform comprehensive theoretical calculations for the optical properties of DAP with different separation distance.
@@
 We find that ON – ONVB DAP is photostable with quantum yields similar to that of the isolated negatively charged ONVB.
-On the contrary, the recombination energy of CB – ONVB develops metastable dim states which act as non-radiative decay pathways.
-The spin flipping in these metastable dim states could mix the doublet and quartet multiplets which causes spin polarization at the $S=1/2$ ground state of the negatively charged ONVB when external magnetic field splits these Kramers-doublets.
+On the contrary, the recombination energy of CB – ONVB develops metastable dim states at certain DAP distances which act as non-radiative decay pathways.
+The spin flipping in these metastable dim states could mix the doublet and quartet multiplets which causes spin-polarization at the $S=1/2$ ground state of the negatively charged ONVB when external magnetic field splits these Kramers-doublets.
 As a consequence, ODMR spectrum Li and Gali (2022) may be observed for the pair of CB – ONVB in the $S=1/2$ ground state in the presence of a constant magnetic field.
 II Results
-Figure 1: (a) Schematics view of CB – ONVB DAP with various distances.
-The numbers labeled here indicate the location of CB.
-(b) The distance-dependent ZPL values.
-(c) The electronic structure of defects in the ground state.
-Number 10 is the isolated ONVB(--) defect for reference.
-ADF-STEM technique is capable of imaging carbon and oxygen atoms near the VB monovacancy, particularly focusing on the innermost nitrogen atoms Li et al.
+Figure 1: CB – ONVB DAPs in hBN.
+a The numbers indicate the location of CB at various distances in hBN lattice (boron and nitrogen atoms are gray and green balls, respectively).
+The CB here is placed to site 1 (brown ball).
+b The distance-dependent ZPL energies in the intrinsically bright optical transition without considering the non-radiative processes.
+c The electronic structure of defects in the ground state.
+Number 10 is the isolated ONVB- defect for reference.
+Our modeling relies on recent ADF-STEM results where they were able to image single and multiple carbon and oxygen atoms adjacent to VB monovacancy, in particular, substituting the innermost nitrogen atoms Li et al.
 (2023).
-However, these monovacancy structures with first-neighbor atom substitutions do not yield the bright emission at 2 eV that is frequently observed in experiments, as discussed in Supplementary Note 1.
-Instead, another interesting phenomenon observed experimentally involves the carbon and oxygen substitutional defects in vacancy-free regions.
-The presence of CB, CN and ON can be readily confirmed, while OB has a high formation energy and therefore is rarely observed Li et al.
+However, the monovacancy structures with multiple first-neighbor atom substitutions do not yield bright ZPL emission at 2 eV that has been frequently observed in experiments, as discussed in Supplementary Note 1.
+Instead, another interesting phenomenon was observed by ADF-STEM where carbon and oxygen substitutional defects exist in vacancy-free regions.
+The presence of CB, CN and ON defects can be readily confirmed, while OB defect has a high formation energy and therefore is rarely observed Li et al.
 (2023).
-A previous theoretical study has indicated the donor characteristics of CB and ON Weston et al.
+A previous theoretical study has indicated the donor characteristics of CB and ON defects Weston et al.
 (2018).
-We emphasize that the oxygen related defect ONVB may exhibit a 2 eV emission in its negative charge state Li and Gali (2022).
+We emphasize that the oxygen related ONVB defect may exhibit 2-eV ZPL emission in its negative charge state Li and Gali (2022).
 Hence, we propose that the observed SPE is ONVB-, with the additional charge supplied by either CB or ON.
@@
 1.
+The atomic sites or ”coordinates” of the donor species with respect to the location of the acceptor species are labeled in Fig.
+1a.
 The first-neighbor site, which is capable of forming a direct bond to oxygen, was not considered due to a potential repulsion interaction between ON and CB Guo et al.
@@
 However, the strong interaction within CB1 and CB2 can change the order of occupied defect levels of ONVB- and the $a_{1}$ level shifts down while $a_{2}$ or $b_{2}$ level shifts up (the presence of CB makes the symmetry $C_{1h}$ so the original $a_{2}$ and $b_{2}$ labels are unified).
-Since the $a_{2}$ ($b_{2}$) orbital has out-of-plane wavefunctions while $b_{1}$ lies in-plane, the transition dipole moment is along out-of-plane direction with very small magnitude at $\sim$0.12 Debye so the optical transition between them is dim.
-The bright emission still comes from the $a_{1}$ to $b_{1}$.
-The large energy separation between $a_{1}$ and $b_{1}$ leads large ZPL, it is 3.7 and 3.9 eV for CB1 and CB2, respectively, and locates at ultraviolet region.
+Since the $a_{2}$ ($b_{2}$) orbital has out-of-plane wavefunctions while $b_{1}$ lies in-plane, the transition dipole moment is along out-of-plane direction with very small magnitude at $\sim$0.12 Debye so the optical transition between them is weak.
+The bright emission still comes from the $a_{1}$ to $b_{1}$ optical transition.
+The large energy separation between $a_{1}$ and $b_{1}$ leads to large ZPL energies.
+It is 3.7 and 3.9 eV for CB1 and CB2, respectively, and locates at ultraviolet region.
 With increased distance, the energy levels’ order restores to the ONVB- character and the wavefunction of CB retrieves $D_{3h}$ symmetry.
@@
 In CB5 case, the empty level from CB is below the $b_{1}$ orbital from ONVB.
-Nevertheless, the transition dipole moment from $a_{1}$ to CB is only 0.16 Debye so we focus on the bright emission localized on ONVB where the ZPL gradually converges to 1.97 eV.
+Nevertheless, the transition dipole moment from $a_{1}$ to CB is only 0.16 Debye so we focus on the bright emission localized on ONVB where the ZPL energies gradually converge to 1.97 eV.
 The above discussed bright emission is a local excitation within ONVB-.
@@
 With Kohn-Sham density functional theory, we could reliably calculate the total energy and geometry of the $S=3/2$ state.
-Supplementary Figure 3 shows the Kohn-Sham energy difference of the $S=1/2$ dim and bright transitions.
-Generally the bright local excitation is independent on the distance between DAP while the dim one shows strong relation.
-The bright transition is the first optical transition in CB3 and CB4 and when CB resides at larger distances from ONVB then the dim transition has lower energy.
-This means there could be a observable non-radiative decay path for the local excited state to decay to the ground state.
-A simple model to capture the dim transition mechanism can be expressed as
+Supplementary Figure 3 shows the Kohn-Sham energy difference for which Kohn-Sham defect states pairs can be associated with the intrinsically dim and bright optical transitions (without considering non-radiative processes and rates here).
+Generally the bright local excitation is independent on the distance between DAP while the dim one shows a strong relation.
+The bright emission is the first optical transition in CB3 and CB4 and when CB resides at larger distances from ONVB then the dim emission has lower energy.
+This means there could be an observable non-radiative decay path for the local excited state to decay to the ground state.
+A simple model to capture the dim emission mechanism can be expressed as
 $E(R_{i})=E_{\text{gap}}-(E_{\text{D}}+E_{\text{A}})+\frac{e^{2}}{\epsilon R_{i}}\text{,}$ (1)
 where $E_{\text{D}}$ and $E_{\text{A}}$ are the energy levels of the donor and the acceptor in the band gap, respectively.
-Here, we use the charge transition level (CTL) of CB($0|+1$) and ONVB($-1|0$) from previous study.
+Here, we use the charge transition level (CTL) of CB($0|+1$) and ONVB($-1|0$) from a previous study Weston et al.
+(2018).
 $\epsilon$ is the dielectric constant of hBN and $R_{i}$ is separated distance between donor and acceptor.
@@
 The total energy of $S=3/2$ state lies with 1.68 eV to 2.36 eV higher than that of the ground state depending on the DAP distance and this means that the bright excited state ($\sim$2.0 eV) and the metastable states can switch their energy position.
-ON – ONVB is different from CB – ONVB case.
-Even with short distance between them, the ONVB- character is still well kept, and this might be due to shallow donor property of ON.
-The empty levels from ON are very close to CBM so they have negligible influence on the optical transition within ONVB.
+ON – ONVB DAPs are different from CB – ONVB DAPs (see Fig.
+2).
+Even with short distance between ON and ONVB, the ONVB- character is still well kept, and this might be due to shallow donor property of ON.
+The empty levels from ON are very close to the conduction band minimum so they have negligible influence on the optical transition within ONVB.
 Except the ON1 site, the ZPL energies are all below 2.23 eV.
 Fig.
-2a shows the transition dipole moment evolution upon the configurations considered.
+3a shows the transition dipole moment evolution upon the configurations considered.
 It converges fast to that of the isolated ONVB- with slightly increased distance between DAP.
 The large energy difference between ON($0|+1$) and ONVB($-1|0$) results in $E(R_{i})$ to be higher than 2.89 eV, e.g., the $S=3/2$ level of ON4 lies at 3.59 eV above the ground state.
-Correspondingly, no metastable states arise for ON+ – ONVB-, which makes the optically and spin active ONVB- counterpart of DAP a photostable emitter.
-Figure 2: (a) Schematics view of ON – ONVB DAP with various distances.
-The numbers labeled here indicate the location of ON.
-(b) The distance-dependent ZPL values.
-(c) The electronic structure of defects in the ground state.
+Correspondingly, no metastable states arise for ON+ – ONVB-, which makes the optically and spin active ONVB- counterpart of DAP a photostable emitter, albeit with a little variation in the ZPL wavelengths depending on the actual distance between the acceptor and the donor.
+Figure 2: ON – ONVB DAPs in hBN.
+a The numbers indicate the location of ON at various distances in hBN lattice (boron and nitrogen atoms are gray and green balls, respectively).
+The ON here is placed to site 1 (red ball).
+b The distance-dependent ZPL energies.
+c The electronic structure of defects in the ground state.
 Number 10 is the isolated ONVB- defect for reference.
-Figure 3: (a) The transition dipole moment evolution upon defect configurations we considered.
+Figure 3: Two types of emission from DAPs.
+(a) The transition dipole moment evolution upon defect configurations we considered is plotted.
 The dash line indicates the transition dipole moment of ONVB-.
-(b)(c) Mechanism of DAP transition in semiconductors.
+The coordinate labels are defined in Figs.
+1 and 2.
+(b) Type 1 and (c) type 2 mechanism of optical transition for DAPs in semiconductors.
+Type 2 may lead to bright emission (see text for details).
 III Discussion
-The fluorescence mechanisms within DAP as shown in Fig.
+The fluorescence mechanisms within DAP as shown in Figs.
 3b,c can be classified in two types according to the relative energy levels of the donor and the acceptor.
-If the donor is lower than the acceptor level then the charge transfer cannot occur spontaneously.
+If the donor is lower than the acceptor level then the charge transfer cannot spontaneously occur.
 Charge transfer can be induced by illumination from $D^{0}$ to $A^{0}$ (neutral charge state) with photon energy $h\nu_{a}$, and this creates ionized donor $D^{+}$ and acceptor $A^{-}$.
@@
 (2019).
-Unlike the above fluorescence mechanisms within DAP, here the photoexcitation is generally active on the orbitals localized on the acceptor, which has a bright optical transition, especially for ON+ – ONVB-.
+Unlike the above fluorescence mechanisms within DAP, here the photoexcitation is generally active on the orbitals localized on the acceptor, which could lead to a bright emission from the excited state.
+In particular, this is the case for ON+ – ONVB-.
 The shallow donor makes the levels of the dim ON0 – ONVB0 states to lie at much higher energies than the bright excited state’s level.
 In CB+ – ONVB- DAP the situation is different, as its optical loop depicted in Fig.
-4.
-The optical excitation first drives the doublet ground state to the bright excited state.
-The small energy difference between bright excited state and the dim metastable states enables the charge transfer from ONVB to CB through non-radiative decay.
+4a.
+The optical excitation first drives the doublet ground state to the optically active doublet excited state.
+The small energy difference between the optically active excited state and the intrinsically dim metastable states enables the charge transfer from ONVB to CB through non-radiative decay.
 The doublet ($D_{s}$) metastable level and the quartet ($Q_{s}$) metastable level are separated by the spin-spin exchange interaction $J$.
-The ground state and the bright doublet excited states are linked to the $D_{s}$ state by either dim optical transition or internal conversion (IC) while they are linked to the $Q_{s}$ by intersystem crossing (ISC) where spin-flip mediates the latter.
+The ground state and the optically active doublet excited states are linked to the $D_{s}$ state by either a weak optical transition or internal conversion (IC) while they are linked to the $Q_{s}$ by intersystem crossing (ISC) where spin-flip mediates the latter.
 By taking the same orbitals in the $D_{s}$ and $Q_{s}$ it is a common wisdom that IC is a much faster process than ISC.
@@
 At zero magnetic field the $D_{s}$ state provides radiative and non-radiative pathways which shorten the optical lifetime of the ONVB- in this DAP when compared to the optical lifetime of ON+ – ONVB- DAP.
+We note that depending on the magnitude of the non-radiative rate ($r1$ rate, see also Supplementary Figure 6) the non-radiative transition towards the metastable $D_{s}$ may dominate in the decay process and then the dim optical transition may occur from $D_{s}$ DAP state.
 At non-zero magnetic fields the Kramers-doublets split in CB+ – ONVB- DAP where the presence of $Q_{s}$ plays an important role in the spin-dependent non-radiative decay.
 For instance, the higher branch $m_{S}=+1/2$ in $D_{s}$ can mix with lower branch of $m_{S}=-3/2$ in $Q_{s}$.
-In other word, the $D_{s}$ acquires some $Q_{s}$ character and vice versa.
+In other words, the $D_{s}$ acquires some $Q_{s}$ character and vice versa.
 Therefore, the $m_{S}=-1/2$ spin sublevel localized on ONVB- in $D_{s}$ can have larger population than $m_{S}=+1/2$ through the non-radiative decay $r3$ (see Fig.
@@
 The $D_{s}$-$Q_{s}$ splitting scales with distance as $R^{-6}$.
-Relatively, the intersystem crossing caused by hyperfine interaction decays slower than the dipolar mixing and spin-orbit-coupling.
-However, this theory might fail with large separation distance between the two defects, e.g., $\sim 5$ nm, when the wavefunction overlap is negligible.
+Relatively, the intersystem crossing caused by hyperfine interaction decays much slower than the dipolar mixing and spin-orbit-coupling.
 For specified distance $R_{i}$, the rate to flip can be evaluated by Fermi Golden rule as
@@
 The distance between the DAP constituents should be long enough to reduce the energy gap $J$ between $D_{s}$ and $Q_{s}$ but not too long to minimize the spin-mixing rate.
-Unfortunately, the direct calculation of $D_{s}$ and $J$ is beyond the scope of the present study because it requires multireference methods that cannot be applied to the supercell size required for the DAP model.
+Unfortunately, the direct calculation of $D_{s}$ and $J$ is beyond the scope of the present study because it requires multi-reference methods that cannot be applied to the supercell size required for the DAP model.
 Nevertheless, an effective Hamiltonian for ONVB0 – CB0 with $S_{1}=1$ and $S_{2}=1/2$ coupled spin pairs at a given magnetic field $B$ may be written as
@@
 (3) as a function of $B$-field (set perpendicular to the hBN sheet) and the $J$ spin exchange interaction as implemented in Easyspin Stoll and Schweiger (2006).
-Figure 4: The optical pumping loop of CB+ ONVB.
-(a) Bright excitation (red) from ground state (GS) to excited state (ES) is localized on ONVB(--).
-Through internal conversion, the ES state can go to doublet IS CB(0) + ONVB(0).
-The doublet IS can transfer to quartet IS through intersystem crossing with energy separation $J>0$.
-The $\pm$ 1/2 and 3/2 spin sublevels are separated by zero field splitting in quartet.
-The black dot indicates spin mixing between doublet and quartet.
-The black circle indicate the population magnitude at ground state.
-(b) The spin sublevel splitting under magnetic field in doublet and quartet.
-The $J$ is --3000, 0 and 3000 MHz, accordingly.
-(c) The population ratio between $|D_{s},+1/2\rangle$ and $|D_{s},-1/2\rangle$ at dim state upon magnetic field.
+Figure 4: The optical pumping loop of CB– ONVB.
+a Bright excitation (red) from ground state (GS) to excited state (ES) is localized on ONVB(--).
+Through internal conversion, the ES state can go to doublet metastable state of CB0 – ONVB0, i.e., $D_{s}$.
+The $D_{s}$ are separated by $J>0$ energy from the quartet metastable state of CB0 – ONVB0, i.e., $Q_{s}$.
+The $\pm$ 1/2 and 3/2 spin sublevels are separated by the zero-field-splitting in $Q_{s}$.
+The black dot indicates spin mixing between $D_{s}$ and $Q_{s}$.
+The black circle indicates the population magnitude in the ground state.
+b The spin sublevels under magnetic field in $D_{s}$ and $Q_{s}$ with $J$ values at $-3000$, 00 and $+3000$ MHz, accordingly.
+c The population ratio between $|D_{s},+1/2\rangle$ and $|D_{s},-1/2\rangle$ in the metastable state at the given magnetic fields.
 Fig.
-4b shows the splitting of spin sublevels under magnetic field for various $J$, which can exhibit different magnitude and sign depending on the relative orientation of coupled spins as appears in our DFT calculation.
-Due to the zero-field splitting of triplet sublevels of ONVB  the $|Q_{s},\pm 1/2\rangle$ always lie low and the main spin mixing is from $|Q_{s},\pm 3/2\rangle$ and $|D_{s},\pm 1/2\rangle$.
+4b shows the splitting of spin sublevels under magnetic field for various $J$ values, which can exhibit different magnitude and sign depending on the relative orientation of coupled spins as appears in our DFT calculation.
+Due to the zero-field splitting of the triplet sublevels of ONVB  the $|Q_{s},\pm 1/2\rangle$ always lie low and the main spin mixing is from $|Q_{s},\pm 3/2\rangle$ and $|D_{s},\pm 1/2\rangle$.
 The mixing is largely due to $|D_{s},-1/2\rangle$ and $|Q_{s},+3/2\rangle$ when $J<0$, while the mixing is between $|D_{s},+1/2\rangle$ and $|Q_{s},-3/2\rangle$ when $J>0$.
-The different coefficient of spin sublevels in eigenstates indicates the distinguishable population and this population difference, or ODMR contrast, could be tuned through magnetic field.
+The different coefficient in the eigenstates of the spin sublevels implies population difference of the spin states which leads to ODMR contrast that could be tuned through constant magnetic fields.
 More specifically, the variation in mixing with the quartet states can result in different non-radiative relaxation rates, leading to hyperpolarization and ODMR contrast.
@@
 It is important to note that for $J=0$, i.e., when two non-interacting defects are present, both doublet states mix equally with the quartet manifold at each magnetic field, resulting in no discernible difference in non-radiative decay.
-To further substantiate the role of spin mixing in mediating the ODMR signal, we calculated the internal conversion rates from the metastable state.
-We approach this process by using the excited state and MS quartet for $r1$, MS quartet and ground state for $r2$, separately.
-With short distance below 1 nm, the small energy difference between bright and dim state and strong electron-phonon coupling lead to fast internal conversion rate around $10^{10}$ MHz.
-Therefore the system relax back to ground state rapidly without detectable optical signal.
-The orbital overlap integral and electron-phonon coupling decay with increasing distance where the $R$ distant dependent value of the integrals may be approximated by the overlap of two Slater orbitals which goes with $e^{-R}$ Silver and Ruedenberg (1968).
-By the two configurations CB5 and CB7, we can extrapolate the electron-phonon coupling and conversion rate to longer distance that cannot be simulated with DFT, as discussed in Supplementary Note 2.
-The estimated conversion rate $r1$ can be 6-8 MHz with 18 Å separation.
-With around 2 nm separation, the present coupled spin pair model could explain the previously reported signature of the ODMR signal from ground state $S=1/2$ defects reported in hBN sheets and nanotubes Scholten et al.
+To further substantiate the role of spin mixing in mediating the ODMR signal, we calculated the internal conversion rates towards and from the MS metastable state.
+We approach this process by using the excited state and MS $D_{s}$ for $r1$, MS $D_{s}$ and doublet ground state for $r2$, separately.
+With short distance below 1 nm, the small energy difference between the optically active doublet excited state and the dim metastable state and the strong electron-phonon coupling lead to fast internal conversion rate around $10^{10}$ MHz.
+Therefore the system relaxes back to ground state rapidly without detectable optical signal.
+The orbital overlap integral and the resulting electron-phonon coupling decay with increasing distance where the $R$ distant dependent value of the integrals may be approximated by the overlap of two Slater orbitals which goes with $e^{-R}$ Silver and Ruedenberg (1968).
+By the two configurations CB5 and CB7, we can extrapolate the electron-phonon coupling and conversion rate to longer $R$ distance between the donor and acceptor that cannot be directly computed with DFT, as discussed in Supplementary Note 2.
+The estimated conversion rate $r1$ is at 6-8 MHz with 18 Å separation which results in competitive radiative and non-radiative pathways from the optically active excited state of the CB– ONVB DAP defects.
+This means that the ZPL wavelength, quantum yield and brightness of these DAP defects depend on the actual distance of the acceptor and donor.
+It is intriguing that both weak radiative and non-radiative ($r3$ internal conversion) decay pathways link the metastable state to the doublet ground state (Supplementary Note 2) that closes the optical spin-polarization loop of these DAP defects.
+With around 2 nm separation between CB and ONVB, the present coupled spin pair model could explain the previously reported signature of ODMR signals from ground state $S=1/2$ defects in hBN sheets and nanotubes Scholten et al.
 (2024); Gao et al.
-(2023c); Guo et al.
+(2023); Guo et al.
 (2023).
-We note that if the metastable state is sufficiently long-living to carry out spin operations (e.g., ST1 defect in diamond Lee et al.
-(2013)) then this spin pair model can be extended to the ODMR centers where the spin-polarization is observed in the metastable state.
+We underline that if the metastable state is sufficiently long-living to carry out spin operations (e.g., ST1 defect in diamond Lee et al.
+(2013)) then this spin-pair model can be extended to the ODMR centers in hBN where the spin-polarization is observed in the metastable state.
 In this case, the ground state could be singlet from an $S=1/2$ acceptor and $S=1/2$ donor whereas the metastable states comprise single and triplet spin pairs.
-We note a recent work (Ref.
-Scholten et al.
-(2024)) during the preparation of our manuscript which observes this phenomenon without adding microscopic models for the underlying processes.
-In summary, we propose the DAP emission is responsible for most of the visible ZPL emission at around 2.0 eV.
+We note recent works (Refs.
+Patel et al.
+(2024); Scholten et al.
+(2024); Robertson et al.
+(2024)) during the preparation of our manuscript which observe this phenomenon without adding microscopic models for the underlying processes.
+In summary, we propose that the DAP emission is responsible for most of the visible ZPL emission at around 2.0 eV.
 The variance in the ZPL wavelengths and optical lifetimes reported in the experiments can be explained by the type and the distance of the donor defect with respect to a key acceptor defect, ONVB.
-Our simulations reveal the ODMR mechanism associated with the spin pairs in certain DAP structures activated in the dim metastable states which is manifested as ODMR detection of the $S=1/2$ ground state at non-zero constant magnetic fields.
+Our simulations reveal the ODMR mechanism associated with the spin pairs in certain DAP structures activated within the dim metastable states which is manifested as ODMR detection of the $S=1/2$ ground state at non-zero constant magnetic fields.
+Our model explains the challenge to produce indistinguishable SPEs in hBN as the location of the defects pairs should be well engineered.
 DAPs in wide band gap semiconductors and materials are quite common structures and are responsible for optical emissions and ODMR signal for Kramers doublet spin systems.
@@
 In this paper, we carry out the density functional theory (DFT) calculations by the Vienna ab initio simulation package (VASP) code Kresse and Furthmüller (1996a, b) with plane wave basis set.
-The valence electrons and core part are described with projector augmented wave (PAW) potentials Blöchl (1994); Kresse and Joubert (1999).
+We applied a plane wave cutoff energy at 450 eV.
+The valence electrons and the core part are described with projector augmented wave (PAW) potentials Blöchl (1994); Kresse and Joubert (1999).
 A $8\times 8$ two-layered supercell model was used to avoid the interactions between periodic images and this is sufficient to use the $\Gamma$-point sampling scheme.
@@
 With mixing parameter $\alpha=0.32$, the hybrid density functional of Heyd, Scuseria, and Ernzerhof (HSE) Heyd et al.
-(2003) could yield experimental optical gap around 6 eV without renormalization energy included and this functional is used to optimize the geometry and calculate electronic structures.
-The convergence threshold for the forces was set to 0.01 eV/Å with cutoff energy at 450 eV.
+(2003) could yield experimental optical gap around 6 eV without electron-phonon renormalization energy included and this functional is used to optimize the geometry and calculate electronic structures.
+The convergence threshold for the forces was set to 0.01 eV/Å.
 $\Delta$SCF method Gali et al.
-(2009) was used to calculate excited states.
-The band alignment is based on the core level energy of one atom far away from the defect center in defective and it in perfect supercell.
+(2009) was used to calculate electronic excited states.
+The band alignment in the charge correction procedure is based on the core level energy of one selected atom far away from the defect center in the defective and the perfect supercell.
+The zero-field-splitting due to dipolar electron-spin electron-spin interaction was calculated within the PAW formalism Bodrog and Gali (2014) as implemented in VASP by Martijn Marsman.
+We applied the spin decontemination method to arrive at the final result Biktagirov et al.
+(2020).
 Author contribution